OpenStudy (kimberly101):

Use equations to solve the following questions David is 5 years younger than his twin brothers. If the sum of their ages is 52, then how old is David?

OpenStudy (kropot72):

Let the age of each twin brother = x years (x-5) + 2x = 52 3x - 5 = 52 3x = 57 x = 57/3

OpenStudy (kropot72):

Can you finish the rest have got the age of each twin?

OpenStudy (ajprincess):

Let the age of David be x. Then the age of one of his twin brothers=x+5 and that will be the age of the othr too. then x+x+5+x+5=52 3x+10=52 3x=52-10 3x=42 x=42/3 x=14

OpenStudy (math_bender_15):

David must be x. Then the age of one of his twin brothers=x+5 and that will be the age of the othr too. then x+x+5+x+5=52 3x+10=52 3x=52-10 3x=42 x=42/3 x=14